OpenStudy (anonymous):

Let f(x)=-2x-1 and g(x)=x^2+2 Find (f o g)(7)

OpenStudy (dmndlife24):

Compose g(7) into the function f, which is what it's asking...to do that, first, replace x with 7 into the function g(x).

OpenStudy (anonymous):

is it -103?

OpenStudy (dmndlife24):

Well we find that g(7) = 51 f(x) = -2x - 1 = -2(51) -1 = -103 So yes, you are correct!

OpenStudy (anonymous):

Thank you.

OpenStudy (dmndlife24):

Yw :)